OpenStudy (anonymous):

f(x)=6x^3+5x^2...is this a function, if so what is the domain and range

OpenStudy (abb0t):

Set the function equal to zero and solve. You can factor out to get two functions \[x^2(6x+5) = 0\] \[x^2 = 0\] and \[6x+5 = 0\]

OpenStudy (zehanz):

You can put any number in and get a new one with your formula. So now you already know two things: yes it's a function and the domain is all real numbers. @abbot: knowing the the zeroes is nice, but of no importance if it's a 3rd degree function. Think about what the range can be: there might be some wiggling going on around the zeroes, but as x grows very large (pos or neg), so does f(x)! You now also know the range...